Account Current Year Prior Year
Net sales (all credit) $445,400 $362,000
Costs of goods sold $220,000 $185,000
Gross profit $225,400 $177,000
Income from operations $72,000 $80,000
Interest expense $8,000 $14,000
Net income $40,000 $35,000
Cash $34,000 $28,000
Accounts receivable, net $42,000 $62,000
Inventory $120,000 $100,000
Prepaid expenses $4,000 $2,000
Total current assets $200,000 $192,000
Total long-term assets $200,000 $230,000
Total current liabilities $100,000 $165,000
Total long-term liabilities $44,000 $95,000
Common stock, no par, $8,000 shares, value $120/share $80,000 $80,000

What was the dollar and % change? ( for net sales, gross profit, net income)

Please show all equations and calculations
Answer #1

Answer

Account Current Year Prior Year Dollar Change - Increase (Decrease) % Increase (Decrease)
[A] [B] [C = A - B] [D = (C/B) x 100]
Net sales (all credit) $445,400 $362,000 $83,400 23.04%
Costs of goods sold $220,000 $185,000 $35,000 18.92%
Gross profit $225,400 $177,000 $48,400 27.34%
Income from operations $72,000 $80,000 ($8,000) -10.00%
Interest expense $8,000 $14,000 ($6,000) -42.86%
Net income $40,000 $35,000 $5,000 14.29%